OpenStudy (ttop0816):

true or false?? sec30°cos60° = 1

OpenStudy (kc_kennylau):

true

OpenStudy (kainui):

secant is the same as just 1/cosine, so you really have: cos(60)/cos(30)=1. So if you multiply both sides by cos(30) you'll get: cos(60)=cos(30) is this true or not? Remember, these are circular functions, so it's possible that even though they have different angles in them, they can still be equal for their outputs.

OpenStudy (haseeb96):

sec30 = 2\sqrt3 cos 60= 1\2 so when we multiply it \[\frac{ 2 }{ \sqrt{3} }\times \frac{ 1 }{ 2 }= 1\sqrt 3 it is not true because left hand side is not equal to right hand side

OpenStudy (kainui):

 remember, on the unit circle cosine tells you the x-value of a triangle made by rotating by that many degrees. So I've drawn the triangles made by going 30 and 60 degrees. Are their x-coordinates the same? If they are, it's true.

OpenStudy (kainui):

hint: @kc_kennylau is wrong this time.